There's 24 mL of saline in 150 mL of pure solution. What is the % of saline in the solution?
Answer by jim_thompson5910(35256)

The percentage of saline is (amount of saline/amount of solution)*100 = (24/150)*100 = 16%


--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------